Sum of two numbers=15
Product of the numbers=36
Find the numbers

Answer by checkley77(12844)

X+Y=15 OR X=15-Y
XY=36
REPLACE X WITH (15-Y) IN THE ABOVE EQUATION & SOLVE FOR Y.
(15-Y)Y=36
15Y-Y^2-36=0
Y^2-15Y+36=0
(Y-12)(Y-3)=0
Y-12=0
Y=12 ANS.
Y-3=0
Y=3 ANS.